what is the measure of the angle if it is 3x+18 and 7x-22? if they are verticle angles?
Answer by mananth(12270)

3x+18 = 7x-22 because vertically opposite angles are equal.
3x-7x=-22-18
-4x=-40
x=-40/-4
x=10
3x+18 = 3*10+18 = 48 deg